OpenStudy (anonymous):

The sum of a number and two-ninths is one and one-third. What is the number? one-ninth negative one-ninth one and one ninth negative one and one ninth

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

n + 2/9 = 1 1/3 n + 2/9 = 4/3 (subtract 2/9 from each side) n = 4/3 - 2/9 (common denominator is 9) n = 12/9 - 2/9 n = 10/9 = 1 1/9